Foreverangel:

I need help please. Determine the difference in salary, in dollars, for an employee who works 52 hours versus one who works 38 hours.

Foreverangel:

1 attachment
surjithayer:

salary who works for 38 hours=10x=10*38=380 $ Salary who works for 52 hours=15(x-40)+400=15(52-40)+400=15*12+400=180+400=580 $

Foreverangel:

Thank you so much

surjithayer:

yw